HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

MILLIMETER-WAVE SPECTROSCOPY OF AMMONIA-WATER WEAKLY BOUNDED COMPLEX

\underline{P. GYAWALI} , R. A. MOTIYENKO, UMR 8523 - PhLAM - Physique des Lasers Atomes et Molécules, Univ. Lille, CNRS, F-59000 Lille, France; L. ZOU, Laboratoire de Physico-Chimie de l’Atmosphère, Université du Littoral Côte

d’Opale, Dunkerque, France; I. KLEINER, Université Paris Cité and Univ. Paris Est Creteil, CNRS, LISA, 75013, Paris, France

The broadband rotational spectra of ammonia-water (NH_{3}-H_{2}O) complex were measured in the frequency range from 50 to 250 GHz using a supersonic-jet emission spectrometer. The NH_{3}-H_{2}O complex exhibits two large amplitude motions (LAMs): almost free internal rotation of ammonia owing to very low torsional barrier (≈ 10 cm^{−1}), and the inversion of water characterized by relatively high barrier (≈ 700 cm^{−1}). In total, about 150 rotational transitions of NH_{3}-H_{2}O were assigned in this study. They were fitted together with the data from previous studies^{1} using the “hybrid” Hamiltonian approach^{2}. The analysis is in progress as we are currently trying to modify the characteristics of supersonic expansion in order to achieve higher rotational temperatures and consequently to measure higher K_{a} transitions. We also present recent modification of the spectrometer that allowed us to improve spectral resolution and observe water inversion tunneling splittings in pure rotational transitions of NH_{3}-H_{2}O. The latest results will be discussed.

This work has been supported by the French PN LEFE and ANR Labex CaPPA through the PIA under Contract No. ANR-11-LABX-0005-01

1P. A. Stockman, R. E. Bumgarner, S. Suzuki, & G. A. Blake, J. Chem. Phys. 96, 2496 (1992); G. T. Fraser & R. D. Suenram, J. Chem. Phys. 96, 7287 (1992)

2I. Kleiner & J. T. Hougen, J. Mol. Spectrosc. 368, 111255 (2020)

p-number: p234  Submitted on Wed, 21 Jun 2023 19:07:27 +0200